OpenStudy (anonymous):

simplify the expression: ((2/b)-(1/3))/((4/b^2)-(1/9))

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[\frac{ 2 }{ b }-\frac{ 1 }{ 3 }\div \frac{ 4 }{ b ^{2} }-\frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

so lets first look at the first one \[\frac{ 2 }{ b}-\frac{ 1 }{ 3 }\] what is the common denominator

OpenStudy (anonymous):

would I have to multiply b to 1/3 and 3 to the 2/b?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

now what about the second one

OpenStudy (anonymous):

so the first one we have \[\frac{ b }{ 3b }-\frac{ 6 }{ 3b}\] right ?

OpenStudy (anonymous):

right, and then the second one would be 36/9b^2 and b^2/9b^2

OpenStudy (anonymous):

right so we have \[\frac{ b -6}{ 3b}\div \frac{ 36-b ^{2} }{ b ^{2} }\] so what do u do now with dividing fractions

OpenStudy (anonymous):

multiply by the recipricol to make it easier?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."